From a circular sheet of radius 4 cm, a circle of radius 3 cm is cut out. Calculate the area of the remaining sheet after the smaller circle is removed.

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The area of the remaining sheet after the smaller circle is removed will be = Area of the entire circle with radius 4 cm – Area of the circle with radius 3 cm

We know,

Area of circle = πr²

So,

Area of the entire circle = π(4)² = 16π cm^2

And,

Area of the circle with radius 3 cm which is cut out = π(3)² = 9π cm^2

Thus, the remaining area = 16π – 9π = 7π cm^2

Thanks